OpenStudy (anonymous):

NEED HELP SO BADLY, FOR AN ASSIGNMENT! in triangle PQR, sinP=(square root 2)/2 and sinQ= (square root 3)/ 2. without calculating the size of any angles, show that the value of sinR= (square root 2)(1+(square root 3)) /4

OpenStudy (anonymous):

is there any more information, mayebe a picture of a triangle?

OpenStudy (anonymous):

no there isnt unfortunately, thank you for replying though.

OpenStudy (anonymous):

its an assignment about compound and double angles if that is any help?

OpenStudy (anonymous):

if sin P = sqrt2/2 it means that P=pi/4 if sin Q = sqrt3/2 ----''------ Q=pi/3 or 2Pi/3 since the sum of the angles of a triangle is pi, there are two possibilitis for R : R=pi-(pi/4+pi/3) and R=pi-(pi/4+2Pi/3)

OpenStudy (anonymous):

thank you!!!!!!!! i cant believe i didnt see that, really means a lot. have a good day :)

OpenStudy (anonymous):

yw